# Stab-QRAM: An All-Clifford Quantum Random Access Memory for Special Data

### v1

@@ Abstract
I Introduction
The field of quantum computing has transitioned from a theoretical framework to an experimental reality, driven by advancements in physical platforms such as superconducting [1] and trapped-ion [2] systems. These developments have enabled increasingly capable quantum processors. However, a key challenge in realizing the potential of transformative quantum algorithms is the efficient encoding of large classical data into quantum states, as this must be done in a way that preserves the computational advantages of quantum algorithms for big data applications. Analogous to Random Access Memory (RAM) in classical computing [3], Quantum Random Access Memory (QRAM) [5, 4, 6] addresses this challenge by performing coherent memory lookups. Specifically, QRAM transforms a superposition of address states $\sum_{i}\alpha_{i}|i\rangle|0\rangle^{\otimes m}$ into a superposition of data states $\sum_{i}\alpha_{i}|i\rangle|f(i)\rangle$ via the unitary operation $U_{f}$. This capability is critical for quantum algorithmic speedups, as it enables simultaneous querying of multiple memory locations in a single coherent operation, providing exponential parallelism in data access unattainable by classical RAM.
@@ I Introduction
To address this gap, we introduce the Stab-QRAM, a novel DS architecture designed for near- and mid-term quantum devices. Our approach makes a strategic trade-off: we sacrifice universality in favor of exceptional resource efficiency by focusing on data with an affine Boolean structure: functions of the form $f(\mathbf{x})=A\mathbf{x}+\mathbf{b}$ over the finite field $\mathbb{F}_{2}$. While the implementability of such functions with Clifford circuits is a known consequence of the Gottesman-Knill theorem [17], our work moves beyond this principle to establish a complete and practical quantum memory framework. We formalize this concept into a complete, reconfigurable QRAM architecture, and provide the first rigorous analysis of its optimal performance. The resulting architecture circumvents the non-Clifford gate bottleneck entirely, achieving $O(\log N)$ space and depth complexity, which makes it uniquely suited for data-intensive tasks.
This paper presents a comprehensive blueprint for the Stab-QRAM. In Section II, we present the theoretical foundations of our model and characterize the Stab-QRAM’s properties. In Section III, we provide a comparative analysis with existing architectures. Finally, in Section IV, we explore promising applications and discuss the potential for efficient implementation of the Stab-QRAM on contemporary quantum hardware platforms.
II Result
@@
Geometric Architecture — The implementation of the affine function $f(x)=Ax+b$ in the Stab-QRAM architecture is modeled by a logical interaction graph $G_{A}=(V,E)$. The vertices $V$ represent the system’s qubits, which are partitioned into two disjoint sets: an $n$-qubit address register $V_{\text{addr}}=\{x_{1},\dots,x_{n}\}$, and an $m$-qubit data register $V_{\text{data}}=\{d_{1},\dots,d_{m}\}$.
@@
By this construction, every edge connects an address qubit to a data qubit, making $G_{A}$ an inherently bipartite graph. This structural property is fundamental to the architecture’s efficient hardware implementation [20] and it lays the foundation for achieving low circuit depth which will be proved later. For example, consider the case where $n=m=5$ with the matrix
$A=\begin{bmatrix}0&1&0&0&1\\ 1&0&0&1&0\\ 0&1&1&0&0\\ 1&1&0&1&0\\ 0&0&1&1&0\end{bmatrix}$.
The corresponding bipartite graph is visualized in Figure 1.
Figure 1: Two visualizations of the Stab-QRAM’s bipartite graph architecture for a specific $5\times 5$ matrix AA. In both layouts, blue circles represent the 5 address qubits ($V_{\text{addr}}$) and red circles represent the 5 data qubits ($V_{\text{data}}$). An edge between an address and a data qubit indicates a required CNOT gate. (a) A standard linear layout. (b) A circular layout to visualize connectivity.
Gate Count and Locality — Gate count serves as a key metric for evaluating circuit resource requirements. Our Stab-QRAM circuit contains no non-Clifford gates, with CNOT gates being the primary operations. The total CNOT count equals the Hamming weight of matrix $A$ (number of non-zero entries). As illustrated in Figure 2, this count scales linearly with matrix density $p$ (probability of $A_{j,k}=1$) and quadratically with dimension $n$ (for $m=n$), matching the expected $E[\text{count}]=m\cdot n\cdot p$. This corresponds to an $O((\log N)^{2})$ scaling, where $N=2^{n}$ represents the data space size, enabling slow growth and strong scalability for large-scale implementations.
Locality is crucial for hardware realization, as the logical bipartite graph $G_{A}$ must map onto physical hardware with limited connectivity. To analyze this, we model the hardware as a $k$-regular graph [21], where each vertex (representing a physical qubit) has exactly $k$ edges connecting it to other vertices. This structure is chosen for its generality, as it abstracts away from specific proprietary architectures while capturing the fundamental connectivity constraints prevalent across various quantum platforms [22, 23]. For concreteness, consider the connectivity in leading platforms: IBM’s 2025 processors, such as the Nighthawk [24], adopt a square-lattice topology with a 4-degree nearest-neighbor connectivity ($k=4$), enabling denser two-qubit gate implementations in a 2D grid-like arrangement. In contrast, Google’s Willow chip [25] features a more irregular tunable coupler architecture with an average connectivity of 3.47, optimized for error-corrected logical qubits but with boundary effects reducing the effective degree for edge qubits. Moreover, by varying the parameter $k$, it facilitates the exploration of different connectivity levels, enabling predictions about performance on future hardware with potentially enhanced connectivity without being limited to current designs.
Circuit Depth — The logical circuit depth, defined as the minimum number of parallel time steps required to execute all gates, is a critical performance metric for quantum algorithms. Our objective is to determine an optimal scheduling strategy for the CNOT and X gates in the Stab-QRAM circuit to minimize this depth. We demonstrate that this scheduling problem can be solved exactly by mapping it to a classic problem in graph theory.
The scheduling problem is elegantly modeled using the bipartite interaction graph $G_{A}=(V,E)$, where vertices represent qubits and edges represent the required CNOT operations. The fundamental physical constraint is that a single qubit can only participate in one CNOT gate per time step. In the graph-theoretic framework, this constraint translates directly to an edge-coloring problem: any two edges (CNOTs) that are incident on the same vertex (qubit) must be assigned to different time layers (colors). Consequently, the minimum circuit depth for the CNOT portion of the circuit is precisely the chromatic index of the graph $G_{A}$—the minimum number of colors needed for a valid edge-coloring.
@@
Figure 4: Analysis of logical circuit depth. (a) For a fixed size ($n=m=50$), the average depth increases with matrix density $p$, remaining well below the theoretical maximum depth of $n+1$ (red dashed line). (b) The average depth scales linearly against the storage capacity $N=2^{n}$ on a logarithmic axis, visually confirming the $O(n)=O(\log N)$ complexity. (c) A scatter plot for randomly generated matrices reveals no direct correlation between the matrix rank over $\mathbb{F}_{2}$ and the circuit depth.
Numerical simulations in Figure 4 underscore Stab-QRAM’s application advantages. Sparse matrices yield minimal depths, suiting sparse-data tasks like quantum machine learning or graph algorithms with reduced latency. Dense matrices, while deeper, maintain robustness below bounds, supporting reliable performance in dense-data scenarios. The $O(\log N)$ scaling enables efficient large-scale quantum memory, outperforming alternatives in expansive computations. The analysis of matrix rank versus depth in Figure 4(c) yields a key insight: there is no direct correlation. This is because rank is a global algebraic property, whereas the circuit depth is dictated by the local graph property of maximum qubit degree, $\Delta$. This distinction is critical for algorithm design, as it shows the Stab-QRAM’s performance hinges on the local distribution of CNOTs rather than the matrix’s global structure. Consequently, sparse matrices, regardless of their rank, are ideal for leveraging the architecture’s inherent parallelism.
III Comparison
@@
The Stab-QRAM occupies a unique and highly advantageous position within the landscape of QRAM designs. As a domain-specific architecture, it leverages the structure of affine Boolean functions to achieve exceptional efficiency, setting it apart from both general-purpose and other domain-specific models.
@@
IV Discussion
In this work, we introduced the Stab-QRAM, a domain-specific architecture designed to overcome a critical bottleneck in quantum computing by enabling highly efficient loading of structured classical data. By deliberately trading the universality of general-purpose designs for exceptional resource efficiency on affine Boolean problems ($f(x)=Ax+b$), Stab-QRAM attains $O(\log N)$ space and depth complexity with zero T-gates. Its Clifford-only construction is inherently compatible with fault-tolerant quantum computing (FTQC) [27], completely eliminating the costly magic state distillation [28] required by non-Clifford QRAMs. This not only slashes the qubit and runtime overheads in FTQC implementations but also lowers the practical threshold for achieving quantum advantage, positioning Stab-QRAM as a powerful enabler for bringing its target applications into the near-future fault-tolerant era.
@@ IV Discussion
Acknowledgment — We thank Frederic T. Chong, Yongshan Ding, Steve Girvin, Allen Mi, Kate Smith and Youtao Zhang for helpful inputs. GL, YG, ZW, and JL are supported in part by the University of Pittsburgh, School of Computing and Information, Department of Computer Science, Pitt Cyber, PQI Community Collaboration Awards, John C. Mascaro Faculty Scholar in Sustainability, NASA under award number 80NSSC25M7057, and Fluor Marine Propulsion LLC (U.S. Naval Nuclear Laboratory) under award number 140449-R08. ZZ acknowledges NSF CAREER Award No. 2317471. This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725.

### v2

@@ Abstract
Introduction — The field of quantum computing has transitioned from a theoretical framework to an experimental reality, driven by advancements in physical platforms such as superconducting [1] and trapped-ion [2] systems. These developments have enabled increasingly capable quantum processors. However, a key challenge in realizing the potential of transformative quantum algorithms is the efficient encoding of large classical data into quantum states, as this must be done in a way that preserves the computational advantages of quantum algorithms for big data applications. Analogous to Random Access Memory (RAM) in classical computing [3], Quantum Random Access Memory (QRAM) [5, 4, 6] addresses this challenge by performing coherent memory lookups. Specifically, QRAM transforms a superposition of address states $\sum_{i}\alpha_{i}|i\rangle|0\rangle^{\otimes m}$ into a superposition of data states $\sum_{i}\alpha_{i}|i\rangle|f(i)\rangle$ via the unitary operation $U_{f}$. This capability is critical for quantum algorithmic speedups, as it enables simultaneous querying of multiple memory locations in a single coherent operation, providing exponential parallelism in data access unattainable by classical RAM.
@@
To address this gap, we introduce the Stab-QRAM, a novel DS architecture designed for near- and mid-term quantum devices. Our approach makes a strategic trade-off: we sacrifice universality in favor of exceptional resource efficiency by focusing on data with an affine Boolean structure: functions of the form $f(\mathbf{x})=A\mathbf{x}+\mathbf{b}$ over the finite field $\mathbb{F}_{2}$. While the implementability of such functions with Clifford circuits is a known consequence of the Gottesman-Knill theorem [17], our work is the first to leverage this principle to construct a complete and practical quantum memory framework. We formalize this concept into a complete, reconfigurable QRAM architecture, and provide the first rigorous analysis of its optimal performance. The resulting architecture circumvents the non-Clifford gate bottleneck entirely, achieving $O(\log N)$ space and depth complexity, which makes it uniquely suited for data-intensive tasks.
This paper presents a comprehensive blueprint for the Stab-QRAM. We begin by establishing the theoretical foundations of our model and characterizing the Stab-QRAM’s properties. This is followed by a comparative analysis with existing architectures to highlight its unique advantages. We conclude by exploring promising applications and discussing the potential for the Stab-QRAM’s efficient implementation on contemporary quantum hardware platforms.
@@
Stab-QRAM’s properties — The implementation of the affine function $f(x)=Ax+b$ in the Stab-QRAM architecture is modeled by a logical interaction graph $G_{A}=(V,E)$. The vertices $V$ represent the system’s qubits, which are partitioned into two disjoint sets: an $n$-qubit address register $V_{\text{addr}}=\{x_{1},\dots,x_{n}\}$, and an $m$-qubit data register $V_{\text{data}}=\{d_{1},\dots,d_{m}\}$.
@@
By this construction, every edge connects an address qubit to a data qubit, making $G_{A}$ an inherently bipartite graph. This structural property is fundamental to the architecture’s efficient hardware implementation [20] and it lays the foundation for achieving $O(\log N)$ optimal logical circuit depth.
The logical circuit depth, defined as the minimum number of parallel time steps required to execute all gates, is a critical performance metric for quantum algorithms. Our objective is to determine an optimal scheduling strategy for the CNOT and X gates in the Stab-QRAM circuit to minimize this depth. We demonstrate that this scheduling problem can be solved exactly by mapping it to a classic problem in graph theory.
The fundamental physical constraint is that a single qubit can only participate in one CNOT gate per time step. In the graph-theoretic framework, this constraint translates directly to an edge-coloring problem: any two edges (CNOTs) that are incident on the same vertex (qubit) must be assigned to different time layers (colors). Consequently, the minimum circuit depth for the CNOT portion of the circuit is precisely the chromatic index of the graph $G_{A}$—the minimum number of colors needed for a valid edge-coloring.
@@
Figure 1: Analysis of logical circuit depth scaling. (a) For a fixed size ($n=m=50$), the average depth increases with matrix density $p$, remaining well below the theoretical maximum. (b) The average depth scales linearly against the storage capacity $N=2^{n}$ on a logarithmic axis, visually confirming the $O(\log N)$ complexity.
Numerical simulations in Figure 1 underscore Stab-QRAM’s application advantages. Sparse matrices yield minimal depths, suiting sparse-data tasks like quantum machine learning or graph algorithms with reduced latency. Dense matrices, while deeper, maintain robustness below bounds, supporting reliable performance in dense-data scenarios. The $O(\log N)$ scaling enables efficient large-scale quantum memory, outperforming alternatives in expansive computations.
Figure 2: Circuit depth versus matrix rank. A scatter plot for randomly generated matrices reveals no direct correlation between the circuit depth and the matrix rank over $\mathbb{F}_{2}$. This highlights that depth is a local property (maximum degree $\Delta$) rather than a global algebraic property.
To further characterize performance, we also analyzed the effect of the matrix rank over $\mathbb{F}_{2}$, a global algebraic property. The results in Figure 2 yield a key insight: there is no direct correlation with circuit depth. This is because rank is a global property, whereas depth is dictated by the local graph property of maximum qubit degree, $\Delta$. This distinction is critical for algorithm design, as it shows the Stab-QRAM’s performance hinges on the local distribution of CNOTs rather than the matrix’s global structure. Consequently, sparse matrices, regardless of their rank, are ideal for leveraging the architecture’s inherent parallelism.
In addition to its optimal depth, another key metric is the resource requirement, quantified by the gate count. Our Stab-QRAM circuit contains no non-Clifford gates, with CNOT gates being the primary operations. The total CNOT count equals the Hamming weight of matrix $A$ (number of non-zero entries). As illustrated in Figure 3, this count scales linearly with matrix density $p$ (probability of $A_{j,k}=1$) and quadratically with dimension $n$ (for $m=n$), matching the expected $E[\text{count}]=m\cdot n\cdot p$. This corresponds to an $O((\log N)^{2})$ scaling, where $N=2^{n}$ represents the data space size, enabling slow growth and strong scalability for large-scale implementations.
Finally, beyond abstract metrics, practical performance hinges on locality—how efficiently the logical bipartite graph $G_{A}$ can be mapped onto physical hardware with limited connectivity. To analyze this, we model the hardware as a $k$-regular graph [22], where each vertex (representing a physical qubit) has exactly $k$ edges connecting it to other vertices. This structure is chosen for its generality, as it abstracts away from specific proprietary architectures while capturing the fundamental connectivity constraints prevalent across various quantum platforms [23, 24]. For concreteness, consider the connectivity in leading platforms: IBM’s 2025 processors, such as the Nighthawk [25], adopt a square-lattice topology with a 4-degree nearest-neighbor connectivity ($k=4$), enabling denser two-qubit gate implementations in a 2D grid-like arrangement. In contrast, Google’s Willow chip [26] features a more irregular tunable coupler architecture with an average connectivity of 3.47, optimized for error-corrected logical qubits but with boundary effects reducing the effective degree for edge qubits. Moreover, by varying the parameter $k$, it facilitates the exploration of different connectivity levels, enabling predictions about performance on future hardware with potentially enhanced connectivity without being limited to current designs.
@@
Comparison — The Stab-QRAM occupies a unique and highly advantageous position within the landscape of QRAM designs. As a domain-specific architecture, it leverages the structure of affine Boolean functions to achieve exceptional efficiency, setting it apart from both general-purpose and other domain-specific models.
@@
Discussion — In this work, we introduced the Stab-QRAM, a domain-specific architecture designed to overcome a critical bottleneck in quantum computing by enabling highly efficient loading of structured classical data. By deliberately trading the universality of general-purpose designs for exceptional resource efficiency on affine Boolean problems ($f(x)=Ax+b$), Stab-QRAM attains $O(\log N)$ space and depth complexity with zero T-gates. Its Clifford-only construction is inherently compatible with fault-tolerant quantum computing (FTQC) [27], completely eliminating the costly magic state distillation [28] required by non-Clifford QRAMs.
@@
Acknowledements — We thank Frederic T. Chong, Yongshan Ding, Steve Girvin, Allen Mi, Kate Smith and Youtao Zhang for helpful inputs. GL, YG, ZW, and JL are supported in part by the University of Pittsburgh, School of Computing and Information, Department of Computer Science, Pitt Cyber, PQI Community Collaboration Awards, John C. Mascaro Faculty Scholar in Sustainability, NASA under award number 80NSSC25M7057, and Fluor Marine Propulsion LLC (U.S. Naval Nuclear Laboratory) under award number 140449-R08. XZ acknowledges the support from the AWS Center for Quantum Computing and the Fu Foundation School of Engineering and Applied Science of Columbia University. ZZ acknowledges NSF CAREER Award No. 2317471. This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725.